OpenStudy (anonymous):

A country's population in 1990 was 156 million. In 1996 it was 162 million. Estimate the population in 2016 using the exponential growth formula. Round your answer to the nearest million.

OpenStudy (anonymous):

ugg hated these problems!!

OpenStudy (anonymous):

I'm ready to poke my eyes out ;)

OpenStudy (anonymous):

i feel ya!! and they take forever!

OpenStudy (wolf1728):

The population increases156,000,000 to 162,000,000 over 6 years 156,000,000 * (1+r)^6 = 162,000,000 (1+r)^6 = 162,000,000 / 156,000,000 (1+r)^6 = 1.03846153846154 6 * log(1+r) = log(1.03846153846154) log(1+r) = 0.01639041618817 / 6 log(1+r) = 0.0027317360313616 1+r = 1.0063098786004 r = .0063098786004 Testing the value of r 156,000,000 * (1+ .0063098786004)^6 = 1996 population 156,000,000 * 1.03846153846154 = 162,000,000 1996 population r = .0063098786004 I think you can use this information to calculate the 2016 population.

OpenStudy (anonymous):

Instructions:Select the correct answer from each drop-down menu. ∆ABC is translated 2 units down and 1 unit to the left. Then it is rotated 90° clockwise about the origin to form ∆A′B′C′. The coordinates of vertex A′ of ∆A′B′C′ are . The coordinates of vertex B′ of ∆A′B′C′ are . The coordinates of vertex C′ of ∆A′B′C′ are .